2. If the total acceleration of an object is 60 m/s2 and its radial acceleration is 30 m/s2 due south, what is tangential acceleration of the object and the direction of the object?

   
If the total acceleration of an object is 60ms2 and its radial acceleration is 30ms2 due south, what is the tangential acceleration of the object? please answer clearly and show steps
Answer #1

Answer:-

The tangential accleration of object is 60 m/s2.and it is denoted by at=60m/s2 .

And its radial acceleration is 30 m/s2 .and it is denoted by ar=30 m/s2   

We have to find tangential acceleration denoted by atan=?

we know at=ar+atan.​​​​​​

thats why, atan=at-ar.

so ,atan=(60 m/s2-30 m/s2)

atan=30 m/s2.

The tangential accelaration is always tangential.

​​​

​​​​​

​​